(8 points) The transition matrix P for a Markov chain is shown below. Find the stable vector W for this Markov chain. Write the entries as whole numbers or fractions in lowest terms T0 0 0 1 P- 1 프 00 1 3 Wー
(6 points) A certain television show airs weekly. Each time it airs, its ratings are categorized as high, average, or low, and transitions between these categorizations each week behave like a Markov chain. If the ratings are high this week, they will be high, average, or low next week with probabilities 80%, 20%, and 0% respectively. If the ratings are average this week, they will be high. average, or low next week with probabilities 10%, 10%, and 80% respectively. If the ratings are low this week, they will be high, average, or low next week with probabilities 80%, 20%, and 0% respectively. Write all answers as integers or decimals. If the ratings are high this week, what is the probability that they will be average two weeks from now? If the ratings are high the first week, what is the probability that they will be average the second week? If the ratings are average the first week, what is the probability that they will be low the third week?
Answer #1

Let W = [a, b, c, d] where a + b + c + d = 1

Then WP = W which gives below set of equations.

(1/4)c + (1/4)d = a

(3/4)c + (3/4)d = b => (1/4)c + (1/4)d = b/3 => a = b/3

b = c

a = d => d = b/3

a + b + c + d = 1

=> (b/3) + b + b + (b/3) = 1

=> 8b/3 = 1

=> b = 3/8

Thus, c = 3/8

d = a = b/3 = 1/8

Thus,

W = [1/8 , 3/8 , 3/8 , 1/8]

Second question,

Let the states H, A and L denote high, average and low ratings.

The transition probability matrix is,

0.8 0.2 0 P0.1 0.1 0.8 0.8 0.2 0

Probability that ratings is average the second week given high this week = P(X2 = A | X0 = H)  

= P(X2 = A , X1 = H, X0 = H) +  P(X2 = A , X1 = A, X0 = H)  

= 0.8 * 0.2 + 0.2 * 0.1 = 0.18

Probability that ratings is average the second week given high the first week = P(X2 = A | X1 = H) = 0.1

Probability that ratings is low the third week given average first week = P(X3 = L | X1 = A)  

= P(X3 = L , X2 = H, X1 = A) +  P(X3 = L , X2 = A, X1 = A)

= 0.1 * 0 + 0.1 * 0.8 = 0.08